OpenStudy (anonymous):

Apply the problem-solving technique of systems of equations to solve the following problem: The sum of two numbers is 50. The difference of the two numbers is 14. What are the two numbers?

OpenStudy (anonymous):

$$ x + y = 50 \land x - y = 14 $$ $$ x = 14 + y \land y = 18 $$ $$ (x,y) = (32, 18) $$

OpenStudy (anonymous):

thanks